# Your Build Scripts Stink: The State of Code Smells in Build Scripts

### v3

@@ Abstract
Build scripts are files that automate the process of compiling source code, managing dependencies, running tests, and packaging software into deployable artifacts. These scripts are ubiquitous in modern software development pipelines for streamlining testing and delivery. While developing build scripts, practitioners may inadvertently introduce code smells. Code smells are recurring patterns of poor coding practices that may lead to build failures or increase risk and technical debt. The goal of this study is to aid practitioners in avoiding code smells in build scripts through an empirical study of build scripts and issues on GitHub. We employed a mixed-methods approach, combining qualitative and quantitative analysis. We conducted a qualitative analysis of 2000 build-script-related GitHub issues. Next, we developed a static analysis tool, Sniffer, to identify code smells in 5882 build scripts of Maven, Gradle, Cmake, and Make files, collected from 4877 open-source GitHub repositories. We identified 13 code smell categories, with a total of 10,895 smell occurrences, where 3184 were in Maven, 1214 in Gradle, 337 in CMake, and 6160 in Makefiles.
Our analysis revealed that Insecure URLs were the most prevalent code smell in Maven build scripts, while Hardcoded Paths/URLs were commonly observed in both Gradle and CMake scripts. Wildcard Usage emerged as the most frequent smell in Makefiles. The co-occurrence analysis revealed strong associations between specific smell pairs of Hardcoded Paths/URLs with Duplicates, and Inconsistent Dependency Management with Empty or Incomplete Tags, indicating potential underlying issues in the build script structure and maintenance practices. Based on our findings, we also recommended strategies to mitigate the existence of code smells in build scripts to improve the efficiency, reliability, and maintainability of software projects.
Index Terms:
Code smells, Build scripts, CI/CD, Software supply chain security, Empirical software engineering
@@ I Introduction
Build scripts are ubiquitous in modern software development, facilitating the automation of complex software application compilation, testing, and deployment processes. Build scripts are used to transform source code into deliverable artifacts. Developers use build tools such as Maven and Gradle to define their respective build systems [1] and maintain consistency and reproducibility across diverse environments [2]. To automate the software development and testing process, the majority of well-maintained software projects make use of build tools like Maven and Gradle [3]. According to JetBrains [4], Maven and Gradle are popular for project building with 71% and 48% usage in the industry, respectively.
As software systems grow in scale and functionality, build scripts become increasingly complex [5] and might require frequent maintenance [6], which can negatively impact overall project quality and developer productivity [7]. Consequently, these scripts become susceptible to various code smells [8].
@@
Prior studies showed that files without code smells exhibit approximately 33%-65% lower risk of fault compared to those containing code smells [17, 18]. Code smells are important predictors of build failures, indicating a link between code quality in build scripts and build reliability [19]. Moreover, broken builds could disrupt team productivity and negatively impact overall project performance [20]. Previous studies have explored the adverse effects of code smells in platforms such as Travis CI [21], Infrastructure as Code (IaC) [22], and source code [23]. However, despite the widespread use of build scripts in modern software development, particularly within Continuous Integration and Continuous Deployment (CI/CD) pipelines, to our knowledge, a systematic analysis of the presence of code smells in the context of build scripts has not been performed yet. Overlooked code smells in build scripts could inadvertently facilitate the spread of bad coding practices, thereby impeding the adaptability and maintainability of software systems [24, 25]. In this paper, we aim to systematically investigate, identify, and quantify the prevalence of code smells in the context of build scripts, employing both static analysis techniques and qualitative assessments.
@@
We address our research questions by examining four types of build scripts: Maven [26], Gradle [27], CMake [28], and Make [29], as well as build script-related GitHub issues, which are collected from open-source GitHub repositories. We conducted a qualitative analysis of 2,000 GitHub issues to identify prevalent code smells, following established guidelines for qualitative research [30]. Next, we collected build scripts from GitHub and leveraged Large Language Models (LLMs) to identify potential code smells in build scripts. We used six general and code-oriented LLMs, i.e., ChatGPT4, ChatGPT4o, Codellama, Llama 8B, Llama 13B, and Mistral, to detect potential code smells in the build script. Through the analysis, we identified a code smell taxonomy and identified 13 code smell categories. Furthermore, we mapped each code smell category to its corresponding Common Weakness Enumeration (CWE) to establish standardized categorizations of the underlying weaknesses. Based on our qualitative study and findings from the language models, we developed a static analysis tool, “Sniffer”, to automatically detect the presence of code smells. Further, we developed an Oracle through a user study to compare our tool’s performance against human evaluation. For our empirical analysis, we gathered 5882 build scripts from 4877 open-source repositories. We measured the occurrence, frequency, smell density, and co-occurrence of the code smells in the selected four types of build scripts. The contributions of our study are outlined below:
@@
    Developed an Oracle to compare our linter performance against the human evaluation of code smells.
@@ II Background and Literature Review
Code smells act as strong indicators of broader software quality issues. Barrak et al. [19] demonstrate that code and test smells are strong predictors of build failures. According to Adams et al. [36], there is a co-evolutionary relationship between source code and build scripts, emphasizing that both need to remain synchronized. Build scripts are prevalently used in automating the compilation, testing, packaging, and deployment of software and are also susceptible to similar code smells. By encapsulating a sequence of commands and configurations, build scripts eliminate the need for manual intervention, reducing the risk of human error and accelerating the software delivery pipeline [8]. However, when smells in build scripts are neglected, they often lead to build failures [37]. The effort of maintaining build systems is nontrivial. Research by McIntosh et al. [5] showed that build maintenance tasks account for up to 27% of the overhead in source code development and 44% in test development. Moreover, 22% of commits and 27% of development tasks directly involve build scripts, highlighting the substantial developer effort spent on managing build infrastructure. Code smells can lead to broken builds and integration delays, disrupting team productivity and software delivery timelines [38]. Zhang et al.[39] analyzed CI performance smells, reporting a 12.4% improvement in build performance when these smells were addressed. Similarly, Saidani and Ouni [21] demonstrated that incorporating bad smell detection enhances build failure prediction accuracy by 4%, reinforcing the value of systematic smell detection.
Despite the importance of build scripts in modern software development workflows, build scripts remain understudied. While source code smells have been extensively studied and cataloged in other codebases [31][40, 41, 42, 43], there is no comprehensive or standardized taxonomy of smells for build scripts that has been studied. Moreover, security incidents such as the XZ Utils backdoor [44] have underscored how weaknesses and negligence in build processes can act as a vector for supply chain attacks. These highlight the gap and need for a systematic analysis of build script quality to identify both maintenance and security-related smells. To address this gap, this study aims to systematically identify and categorize code smells specific to build scripts across multiple build systems and to suggest their potential mitigation to enhance the maintainability and security of build automation processes.
@@ III-A1 Qualitative Analysis on GitHub Issues
Data Collection: We performed a keyword-based search within the GitHub repositories using the Rest API and GraphQL. We specifically searched for build scripts and code-smell-related keywords in the GitHub repositories and issues’ metadata, including titles, descriptions, and labels for the selected build type. Two authors collaboratively reviewed and iteratively refined the search queries for their relevance and accuracy based on prior studies on code smells and practitioner blog posts [48, 34, 49, 50, 51]. The search strings included: “codesmell”, “code smell”, “bad practice”, “bad smell”, “HTTP external download”, “empty password”, “long parameter”, “long method”, “inconsistent version”, “abandoned dependencies”, “unused dependency”, “duplicate code”, “long method”, “admin by default”, “anti-pattern”, “hardcoded paths”, “duplicated configuration”, “duplicate code”, “conditional complexity”, “Dead code”, “inconsistent name”, “complex build logic”, “inconsistent dependency versions”, “unused dependencies”, “excessive build times”, “lack of modularity”, “excessive task dependencies”, “Path-misconfiguration”. Following the process, we collected 7533 GitHub issues. Further removing the duplicates, 7104 unique issues remained.
@@
Qualitative Coding: We conducted a qualitative analysis to obtain a summarized overview of recurring bad coding practices in build scripts and to gain insight into potential code smells. For this process, we randomly selected 200 issues of each type of build script (Maven, Gradle, CMake, and Make), totaling 800 GitHub issues out of 7104 issues. Two authors independently conducted a descriptive qualitative analysis [52] on these sampled issues. To our knowledge, at the time of this study, no prior research or predefined list of code smells specific to build scripts was available. To enhance the completeness of our qualitative findings, we compared the 800 issues analyzed with the remaining dataset using the Jaccard similarity index [53]. We employed the Jaccard score to identify and prioritize dissimilar issues for further analysis based on the assumption that similar issues are likely to yield similar code smell detections as in the first 800 issues. Consequently, issues with the lowest Jaccard scores were selected as starting points for further investigation. We continued iterative analysis of GitHub issues until no new code smells emerged. Following the process, we analyzed an additional 1200 issues, 300 issues from each building type, ultimately reaching a saturation point [54], where no new code smells were identified. Following the process, we qualitatively coded 2000 GitHub issues. After conducting descriptive coding, the authors collaboratively finalized the code smells list and addressed any conflicts through a negotiated agreement approach [55]. Disagreements were resolved either by eliminating inappropriate categories for code smells or by merging closely related categories to form a unified and coherent categorization. Following the processes, we made an initial list of code smells based on GitHub issue analysis [56].
@@ III-A2 Qualitative Analysis of Build Scripts Leveraging LLMs
Validation of LLM: To evaluate the models’ performance, three authors independently and manually analyzed the same 100 scripts to identify code smells. The results from this manual analysis were then compared against the results provided by the six LLMs’ outputs. To avoid potential bias, the first author was responsible for executing the LLMs and comparing their output with the manual analysis, while the second, third, and fourth authors conducted the manual evaluations and resolved any conflict through iterative discussion and following negotiation agreement [55]. Among the analyzed LLMs, ChatGPT-4o showed the highest recall, accurately identifying the majority of true positive code smells, presented in Table I. As a result, ChatGPT-4o was selected for further analysis.
Code Smell Taxonomy: We analyzed 2,134 collected build scripts using ChatGPT-4o, following a consistent prompt and procedure. This process provided a large list of code smells. To facilitate structured analysis, we grouped similar code smells into broader smell categories. For instance, LLM-generated responses such as “HTTP URL for Maven Central,” “Insecure Repository URL,” and “HTTP Repository URL” were grouped under the category “Insecure URLs,” as they reflect a common pattern of using unsecured links within build scripts. We merged findings from our GitHub repository analysis with the LLM-generated results, ensuring that the final taxonomy was grounded in both automated detection and manual findings. Two coders independently conducted the analysis, with iterative refinement until thematic saturation was reached [54]. Through this process, we identified 13 distinct code smell categories. The inter-rater reliability for categorization was substantial, with a Cohen’s k of 0.81. Our achieved inter-rater reliability score indicates strong agreement between coders. The final and complete list of identified categories indicating their existence in different build types is presented in Table II.
@@ III-A3 Code Smells to CWE Mapping
To assess the security relevance of the identified code smells in build scripts, we mapped each instance to corresponding entries in the Common Weakness Enumeration (CWE) database [60]. We selected CWE as the reference framework due to its standardized taxonomy of software security weaknesses and proper maintenance by the cybersecurity community. For example, the smell “Hardcoded credential” was mapped to both CWE-798: Use of Hard-coded Credentials and CWE-259: Use of Hardcoded Password [60]. To be consistent with the objectivity in the mapping process, two authors independently performed mapping between code smells and CWEs. Further, we resolved our conflicts through discussion with another author. Our evaluations demonstrate perfect consistency, resulting in a Cohen’s Kappa score of 1.0, which indicates perfect inter-rater reliability. A list of the identified code smells across different types of build scripts with the associated CWE is in Table II.
@@ III-B Answer to RQ1: What code smells occur in build scripts?
  • Complexity: This code smells refers to overly complex logic in build scripts such as nested conditionals, inline shell commands, or convoluted plugin chains. This smell is also known as conditional complexity, using lengthy, cascading if statements or switch/case [63]. The existence of such smells reduces maintainability and increases the likelihood of misconfiguration. This smell is linked to CWE-710: Improper Adherence to Coding Standards.
@@
  • Hardcoded Credentials: Hardcoded credentials refer to the recurring pattern of embedding authentication or sensitive information, such as usernames, passwords, API tokens, or private keys, directly within build scripts. This approach poses a major security risk as sensitive information may be inadvertently exposed through version control. As a result, secrets can be exposed to public repositories and cause unauthorized access to the system [22]. Prior work shows this smell in IaC and CI/CD (e.g., secrets in Puppet, workflow tokens), with prolonged lifetimes and high exploitation risk. In build scripts, the risk is amplified because credentials are embedded in artifact management and dependency resolution workflows, directly affecting supply-chain integrity. This smell aligns with CWE-798: Use of Hard-coded Credentials and CWE-259: Use of Hard-coded Password, which highlights the vulnerability of storing sensitive data in an unprotected and immutable manner.

### v4

@@ Abstract
Build scripts automate the process of compiling source code, managing dependencies, running tests, and packaging software into deployable artifacts. These scripts are ubiquitous in modern software development pipelines for streamlining testing and delivery. While developing build scripts, practitioners may inadvertently introduce code smells, which are recurring patterns of poor coding practices that may lead to build failures or increase risk and technical debt. The goal of this study is to aid practitioners in avoiding code smells in build scripts through an empirical study of build scripts and issues on GitHub. We employed a mixed-methods approach, combining qualitative and quantitative analysis. First, we conducted a qualitative analysis of 2000 build-script-related GitHub issues to understand recurring smells. Next, we developed a static analysis tool, Sniffer, to automatically detect code smells in 5882 build scripts of Maven, Gradle, CMake, and Make files, collected from 4877 open-source GitHub repositories. To assess Sniffer’s performance, we conducted a user study, where Sniffer achieved higher precision, recall, and F-score. We identified 13 code smell categories, with a total of 10,895 smell occurrences, where 3184 were in Maven, 1214 in Gradle, 337 in CMake, and 6160 in Makefiles.
Our analysis revealed that Insecure URLs were the most prevalent code smell in Maven build scripts, while Hardcoded Paths/URLs were commonly observed in both Gradle and CMake scripts. Wildcard Usage emerged as the most frequent smell in Makefiles. The co-occurrence analysis revealed strong associations between specific smell pairs of Hardcoded Paths/URLs with Duplicates, and Inconsistent Dependency Management with Empty or Incomplete Tags, which indicate potential underlying issues in the build script structure and maintenance practices. Based on our findings, we also recommended strategies to remove code smells in build scripts to improve the efficiency, reliability, and maintainability of software projects.
@@ I Introduction
Build scripts are ubiquitous in modern software development, facilitating the automation of complex software application compilation, testing, and deployment processes. Build scripts are used to transform source code into deliverable artifacts. Developers use build tools such as Maven and Gradle to define their respective build systems [1] and maintain consistency and reproducibility across diverse environments by standardizing the build process [2, 3, 4]. To automate the software development and testing process, the majority of well-maintained software projects make use of build tools like Maven and Gradle [5]. According to JetBrains [6], Maven and Gradle are popular for project building with 71% and 48% usage in the industry, respectively.
As software systems grow in scale and functionality, build scripts become increasingly complex [7] and might require frequent maintenance [8], which can negatively impact overall project quality and developer productivity [9]. Consequently, these scripts become susceptible to various code smells.
@@
Prior studies showed that files without code smells exhibit approximately 33%–65% lower risk of fault compared to those containing code smells [18, 19]. Code smells are important predictors of build failures, indicating a link between code quality in build scripts and build reliability [20]. Moreover, broken builds could disrupt team productivity and negatively impact overall project performance [21]. Previous studies have explored the adverse effects of code smells in platforms such as Travis CI [22], Infrastructure as Code (IaC) [23], and source code [24]. However, despite the widespread use of build scripts in modern software development, particularly within Continuous Integration and Continuous Deployment (CI/CD) pipelines, to our knowledge, a systematic analysis of the presence of code smells in the context of build scripts has not been performed yet. Overlooked code smells in build scripts could inadvertently facilitate the spread of bad coding practices, thereby impeding the adaptability and maintainability of software systems [25, 26]. In this paper, we aim to systematically investigate, identify, and quantify the prevalence of code smells in the context of build scripts, employing both static analysis techniques and qualitative assessments.
@@
We address our research questions by examining four types of build scripts: Maven [27], Gradle [28], CMake [29], and Make [30], as well as build script-related GitHub issues, which are collected from open-source GitHub repositories. We conducted a qualitative analysis of 2,000 GitHub issues to identify prevalent code smells, following established guidelines for qualitative research [31]. Next, we collected build scripts from GitHub and leveraged Large Language Models (LLMs) to identify potential code smells in build scripts. We used six general and code-oriented LLMs, i.e., ChatGPT4, ChatGPT4o, Codellama, Llama 8B, Llama 13B, and Mistral, to detect potential code smells in the build script. Through the analysis, we identified a code smell taxonomy and identified 13 code smell categories. Furthermore, we mapped each code smell category to its corresponding Common Weakness Enumeration (CWE) to establish standardized categorizations of the underlying weaknesses. Based on our qualitative study and findings from the LLMs, we developed a static analysis tool, “Sniffer”, to automatically detect code smells. Further, we developed an Oracle through a user study to compare our tool’s performance against human evaluation. For our empirical analysis, we gathered 5882 build scripts from 4877 open-source repositories. We measured the occurrence, frequency, smell density, and co-occurrence of the code smells. The contributions of our study are outlined below:
@@
    Development of an Oracle to compare our linter performance against the human evaluation of code smells.
@@ II Background and Literature Review
Code smells act as strong indicators of broader software quality issues. Barrak et al. [20] demonstrate that code and test smells are strong predictors of build failures. According to Adams et al. [37], there is a co-evolutionary relationship between source code and build scripts, emphasizing that both need to remain synchronized. Build scripts are prevalently used in automating the compilation, testing, packaging, and deployment of software and are also susceptible to similar code smells. By encapsulating a sequence of commands and configurations, build scripts eliminate the need for manual intervention, reducing the risk of human error and accelerating the software delivery pipeline [38]. However, when smells in build scripts are neglected, they often lead to build failures [39]. The effort of maintaining build systems is nontrivial. Research by McIntosh et al. [7] found that build maintenance tasks account for up to 27% of the overhead in source code development and 44% in test development. Moreover, 22% of commits and 27% of development tasks directly involve build scripts, highlighting the substantial developer effort spent on managing build infrastructure. Code smells can lead to broken builds and integration delays, disrupting team productivity and software delivery timelines [40]. Zhang et al. [41] analyzed CI performance smells, reporting a 12.4% improvement in build performance when these smells were addressed. Similarly, Saidani and Ouni [22] demonstrated that incorporating bad smell detection enhances build failure prediction accuracy by 4%, reinforcing the value of systematic smell detection.
Despite the importance of build scripts in modern software development workflows, build scripts remain understudied. While source code smells have been extensively studied and cataloged in other codebases [32, 42, 43, 44, 45], there is no comprehensive or standardized taxonomy of smells for build scripts that has been studied. Moreover, security incidents such as the XZ Utils backdoor [46] have underscored how weaknesses and negligence in build processes can act as a vector for supply chain attacks. These highlight the gap and need for a systematic analysis of build script quality to identify both maintenance and security-related smells. To address this gap, this work aims to systematically identify and categorize code smells specific to build scripts across multiple build systems and to suggest their potential mitigation to enhance the maintainability and security of build automation processes.
@@ III-A1 Qualitative Analysis on GitHub Issues
Data Collection: We performed a keyword-based search within the GitHub repositories using the Rest API and GraphQL. We specifically searched for build scripts and code-smell-related keywords in the GitHub repositories and issues’ metadata, including titles, descriptions, and labels for the selected build type. Two authors collaboratively reviewed and iteratively refined the search queries for their relevance and accuracy based on prior studies on code smells and practitioner blog posts [50, 35, 51, 52, 53]. The search strings included:
“codesmell”, “code smell”, “bad practice”, “bad smell”, “HTTP external download”, “empty password”, “long parameter”, “long method”, “inconsistent version”, “abandoned dependencies”, “unused dependency”, “duplicate code”, “long method”, “admin by default”, “anti-pattern”, “hardcoded paths”, “duplicated configuration”, “duplicate code”, “conditional complexity”, “Dead code”, “inconsistent name”, “complex build logic”, “inconsistent dependency versions”, “unused dependencies”, “excessive build times”, “lack of modularity”, “excessive task dependencies”, “Path-misconfiguration”.
Following the process, we collected 7533 GitHub issues. Further removing the duplicates, 7104 unique issues remained.
@@
Qualitative Coding: We conducted a qualitative analysis to obtain a summarized overview of recurring bad coding practices in build scripts and to gain insight into potential code smells. For this process, we randomly selected 200 issues of each type of build script (Maven, Gradle, CMake, and Make), totaling 800 GitHub issues out of 7104 unique issues. Two authors independently conducted a descriptive qualitative analysis [54] on these sampled issues. To our knowledge, at the time of this study, no prior research or predefined list of code smells specific to build scripts was available. To enhance the completeness of our qualitative findings, we compared the 800 issues analyzed with the remaining dataset using the Jaccard similarity index [55]. We employed the Jaccard score to identify and prioritize dissimilar issues for further analysis based on the assumption that similar issues are likely to yield similar or the same code smell as detected in the first 800 issues. Consequently, issues with the lowest Jaccard scores were selected as starting points for further investigation. Following the process, we analyzed an additional 1200 issues, 300 issues from each building type, ultimately reaching a saturation point [56], where no new code smells were identified. Following this process, we qualitatively coded 2000 GitHub issues in total. After conducting descriptive coding, the authors collaboratively finalized the code smells list and addressed any conflicts through a negotiated agreement approach [57]. Disagreements were resolved either by eliminating inappropriate categories for code smells or by merging closely related categories to form a unified and coherent categorization. Following the processes, we created an initial list of code smells based on our GitHub issue analysis [58].
@@ III-A2 Qualitative Analysis of Build Scripts Leveraging LLMs
Validation of LLM: To evaluate the models’ performance, three authors independently and manually analyzed a selection of 100 build scripts to verify detected code smells. The results from this manual analysis were then compared against the results provided by the six LLMs’ outputs. To avoid potential bias, the first author was responsible for executing the LLMs and comparing their output with the manual analysis, while the second, third, and fourth authors conducted the manual evaluations and resolved any conflict through iterative discussion and following negotiation agreement [57]. Among the analyzed LLMs, ChatGPT-4o showed the highest recall, accurately identifying the majority of true positive code smells, presented in Table I. As a result, ChatGPT-4o was selected for further analysis.
Code Smell Taxonomy: We analyzed 2,134 collected build scripts using ChatGPT-4o, following a consistent prompt and procedure. This process provided a large list of code smells. To facilitate structured analysis, we grouped similar code smells into broader smell categories. For instance, LLM-generated responses such as “HTTP URL for Maven Central,” “Insecure Repository URL,” and “HTTP Repository URL” were grouped under the category “Insecure URLs,” as they reflect a common pattern of using unsecured links within build scripts. We merged findings from our GitHub repository analysis with the LLM-generated results, ensuring that the final taxonomy was grounded in both automated detection and manual findings. Two coders independently conducted the analysis, with iterative refinement until thematic saturation was reached [56]. Through this process, we identified 13 distinct code smell categories. The inter-rater reliability for categorization was a Cohen’s k of 0.81, which indicates strong agreement between coders. The final and complete list of identified categories is provided in Table II.
@@ III-A3 Code Smells to CWE Mapping
To assess the security relevance of the identified code smells in build scripts, we mapped each instance to corresponding entries in the Common Weakness Enumeration (CWE) database [62]. We selected CWE as the reference framework due to its standardized taxonomy of software security weaknesses and proper maintenance by the cybersecurity community. For example, the smell “Hardcoded credential” was mapped to both CWE-798: Use of Hard-coded Credentials and CWE-259: Use of Hardcoded Password [62]. To be consistent with the objectivity in the mapping process, two authors independently performed mapping between code smells and CWEs. Further, we resolved our conflicts through discussion with another author. Our evaluations demonstrate perfect consistency, resulting in a Cohen’s Kappa score of 1.0 after resolving conflict, which indicates perfect inter-rater reliability. A list of the identified code smells across different types of build scripts with the associated CWE is in Table II.
@@ III-B Answer to RQ1: What code smells occur in build scripts?
  • Complexity: This code smell refers to overly complex logic in build scripts, such as nested conditionals, inline shell commands, or convoluted plugin chains. This smell is also known as conditional complexity, using lengthy, cascading if statements or switch/case [65]. The existence of such smells reduces maintainability and increases the likelihood of misconfiguration. This smell is linked to CWE-710: Improper Adherence to Coding Standards.
@@
  • Hardcoded Credentials: Hardcoded credentials refer to the recurring pattern of embedding authentication or sensitive information, such as usernames, passwords, API tokens, or private keys, directly within build scripts. This approach poses a major security risk as sensitive information may be inadvertently exposed through version control. As a result, secrets can be exposed to public repositories and cause unauthorized access to the system. Prior work shows this smell in IaC and CI/CD (e.g., secrets in Puppet, workflow tokens), with prolonged lifetimes and high exploitation risk [23, 64]. In build scripts, the risk is amplified because credentials are embedded in artifact management and dependency resolution workflows, directly affecting supply-chain integrity. This smell aligns with CWE-798: Use of Hard-coded Credentials and CWE-259: Use of Hard-coded Password, which highlights the vulnerability of storing sensitive data in an unprotected and immutable manner.